# Junhao Chen* 1122    Xiang Li* 33    Xiaojun Ye 1144    Chao Li 55 Zhaoxin Fan 66

### v1

@@ 3.4 Draft 3D Model Selection
Then the LMM agent 5 for model selection in Idea-2-3D selects the superior draft 3D model $D_{\rm best}$ from the generated set $\{D_{0},D_{1},...,D_{N-1}\}$ based on the fidelity and relevance to the user’s IDEA. This critical step filters out subpar models, ensuring high-quality iterative generations.
@@
Here, $p_{\rm select}$ is the prompt for the Large Multimodal Model (LMM), guiding the selection of the best draft 3D model. This mechanism compensates for the discrepancy observed when high-quality T-2-I outputs do not necessarily translate into satisfactory I-2-3D models. By assessing the semantic coherence and visual quality of N similar draft models, the LMM identifies the best one. This comparative analysis, akin to a find the difference task, is impossible to achieve by conventional techniques, yet state-of-the-art LMMs like GPV-4V [29, 26] have demonstrated reliable performance in this selection process.
@@ 3.7 Memory Module
The memory module $m$ is integral to the Idea-2-3D framework, serving as a repository for data accrued over the iterative process, as shown by the blue rectangle in Fig. 3. It stores feedback, selected draft 3D models, and corresponding text prompts in a structured 3D model-image-text sequence, enabling the LMM to leverage past experiences and insights gained during previous iterations.
@@
In this representation, $P_{\rm iter}^{*},D_{\rm iter}^{*},F_{\rm iter}$ represent the optimal text prompt, 3D model, and textual feedback from each $\rm iter^{th}$ iteration. The memory module $m$ aids the agent in identifying specific T-2-3D output traits, such as misunderstood keywords. This knowledge is then integrated into generating refined 3D model prompts, enhancing the precision and adaptability of model generation. If T-2-3D struggles with certain design aspects, $m$ guides subsequent iterations to optimize prompts more effectively, ensuring continuous improvement in Idea-2-3D, thereby increasing its alignment with complex IDEAs.
@@ 4.1 Experimental Setup
Now we describe our experiment settings. Recall that our method exploits LMM-based agents and off-the-shelf T-2-I and I-2-3D models. As shown in Table. 1, our LMM agent is by default based upon GPT4-V if not specified but we also present results using LLaVA since GPT4-V is not open-sourced. We also experiment with different T-2-I and I-2-3D models to provide as many data points as possible.
@@
T-2-3D baseline. Our first baseline is caption-based. Since no former 3D AIGC methods can be used for IDEA input, we convert image inputs and 3D model inputs (after multiview rendering) into textual descriptions by captioning them with the LMM. In Fig. 2 and Table. 1, this baseline is called T-2-3D.
@@
Ours w/o iterative refinement. To demonstrate the impact of an iterative self-refinement design, we construct an Idea-2-3D variant with only one iteration. Compared with caption-based baselines, this one features multiple prompt generation and best 3D model selection. This baseline is noted in Table. 1 as Idea-2-3D in the Initial-round column.
Lastly, the full Idea-2-3D exploits by a default cap of 5 iterations if not specified, noted in Table. 1 as Idea-2-3D in the Iterative self-refined column.
4.2 Evaluation Protocol
Following the evaluation practice of Parti [56], we provide a set of prompts for evaluation. Our prompt set involves 20 diverse user-generated IDEAs, including examples from Fig. 1 and Fig. 2. These IDEAs span a range of complexities: nine were text-only, five featured a single image or 3D model, and six included two images or 3D models as multimodal inputs. Each test case was meticulously designed to represent real-world scenarios.
The primary metric for our evaluation was a user preference study. Participants were presented with various IDEAs alongside multiple 3D model outputs from both caption-based T-2-3D baselines and Idea-2-3D. To aid decision-making, users viewed rotating video representations of each 3D model. The order of presentation for both the cases and the model outputs was randomized to avoid any sequence bias, ensuring an unbiased assessment of the users’ preferences.
For Table. 1, we evaluate two T-2-I mdoels: DeepFloyd IF [14], and DALL·E [27, 28], two I-2-3D models: Zero123 [9, 20] and Wonder3D [22], and two LMM agent options: GPT-4V [29, 26] and LLaVA-v1.5-13b [19, 18].
Figure 4: Comparison between prompts for "Caption-based T-2-I prompt for initial round", "Idea2-3-D prompt for initial round" and "Iterative self-refinement Idea-2-3D end round" and the comparison between generated 3D models.
Figure 5: An example of the process for a 3D model generated through the Idea-2-3D framework. An initial prompt is first generated based on the input IDEA and $p_{\text{gen}}$, then multiple images are generated from the T-2-I model, and then the corresponding draft 3D model is generated for each image separately. The most appropriate 3D model is selected based on $p_{\text{select}}$ and feedback is generated. The results are output after several rounds of iterations. Green rectangles indicate text prompts for T-2-3D. blue rectangles indicate: modification suggestions given by agent based on draft 3D models and $p_{\text{fb}}$.
4.3 Results
4.3.1 Quantitative user study.
@@ 4.3 Results
Our evaluation, detailed in Tab. 1, compares caption-based T-2-3D baselines and Idea-2-3D using models specified in each row. We ask users to evaluate which model (the T-2-3D model, the first round results of Idea-2-3D, and the end round results of Idea-2-3D) is more satifying, as well as evaluating each 3D model for compliance with IDEA.
Table 1: This user study’s comparative analysis reveals the remarkable superiority of the Idea-2-3D framework over existing caption-based T-2-3D methods.
@@
Comparison stage: T-2-3D vs. Idea-2-3D (Initial-round) Idea-2-3D (Iterative self-refined)
@@
Satisfying
IDEA
T-2-I I-2-3D LMM T-2-3D Idea-2-3D T-2-3D Idea-2-3D T-2-3D Idea-2-3D T-2-3D Idea-2-3D
DALL·E Zero123 GPT4-V 25.8% 74.2% 41.5% 78.2% 6.5% 93.5% 41.5% 94.2%
DALL·E Zero123 LLaVA 27.5% 72.5% 70.6% 29.7% 70.3% 82.3%
DALL·E Wonder3D GPT4-V 18.8% 81.2% 47.9% 74.3% 3.5% 96.5% 47.9% 91.1%
DALL·E Wonder3D LLaVA 16.3% 83.7% 69.4% 10.3% 89.7% 76.2%
DeepFloyd IF Zero123 GPT4-V 20.0% 80.0% 38.5% 64.1% 28.3% 71.7% 38.5% 73.7%
DeepFloyd IF Zero123 LLaVA 29.0% 71.0% 57.6% 34.0% 66.0% 66.8%
DeepFloyd IF Wonder3D GPT4-V 29.3% 70.7% 32.6% 66.1% 14.8% 85.2% 32.6% 75.9%
DeepFloyd IF Wonder3D LLaVA 26.0% 74.0% 55.3% 18.3% 81.7% 64.5%
@@
The results show that Idea-2-3D markedly enhances user preference scores across a diverse range of T-2-I, I-2-3D, and LMM models. Notably, the Idea-2-3D framework’s initial prompting stage significantly outperforms the caption-based T-2-3D results by effectively decomposing and interpreting the user’s multimodal IDEA, thereby selecting the most suitable 3D model. This improvement is further amplified in the iterative self-refinement stage of Idea-2-3D.
@@
For instance, in scenarios utilizing DALL-E, Zero123, and GPT-4V, Idea-2-3D models were preferred by users in 74.2% of the cases over T-2-3D in the initial round, demonstrating a higher IDEA satisfaction rate (78.2%). Conversely, T-2-3D models achieved only 41.5% satisfaction. In the iterative refinement comparisons, Idea-2-3D models were favored even more (93.5% preference), achieving a remarkable 94.2% satisfaction rate.
@@
These findings were consistent across various T-2-I, I-2-3D, and LMM configurations. Additionally, we observed that stronger T-2-I models, with enhanced language understanding capabilities, contributed to improved performance in Idea-2-3D, suggesting that our framework may enjoy the development of off-the-shelf models it invokes.
4.3.2 Visualization of the effects of the self-iterative refinement process.

### v2

@@ 3.4 Draft 3D Model Selection
Then the LMM agent 5 for model selection in Idea23D selects the superior draft 3D model $D_{\rm best}$ from the generated set $\{D_{0},D_{1},...,D_{N-1}\}$ based on the fidelity and relevance to the user’s IDEA. This critical step filters out subpar models, ensuring high-quality iterative generations.
@@
Here, $p_{\rm select}$ is the prompt for the Large Multimodal Model (LMM), guiding the selection of the best draft 3D model. $f_{select}$ uses specific few-shot prompts for the LMM. It renders six views of each 3D model, combines them into a single image, and then inputs this image into the LMM. The LMM then selects the 3D model that most closely matches the user’s IDEA input to serve as the draft model for the current iteration. This mechanism compensates for the discrepancy observed when high-quality T-2-I outputs do not necessarily translate into satisfactory I-2-3D models. By assessing the semantic coherence and visual quality of N similar draft models, the LMM identifies the best one. This comparative analysis, akin to a find the difference task, is impossible to achieve by conventional techniques, yet state-of-the-art LMMs like GPT-4V OpenAI (2023a), GPT-4o OpenAI (2024) and InternVL Chen et al. (2024b) have demonstrated reliable performance in this selection process.
@@ 3.6 Revised Prompt Generation
Figure 3: Overview of 3D models generated from various types of multimodal IDEA inputs supported by Idea23D. The light red box on the left is the user input IDEA containing text, images and 3D models. In the center are the baseline results generated directly from the same T-2-I model with caption-based T-2-I prompt (see Sec. 4.1). The model on the right is the result generated by iteratively self-refining the T-2-I prompts with Idea23D. Comparison with more existing methods is shown in Fig. 7.
@@ 3.7 Memory Module
The memory module $m$ is integral to the Idea23D framework, serving as a repository for data accrued over the iterative process, as shown by the blue rectangle in Fig. 2. It stores feedback, selected draft 3D models, and corresponding text prompts in a structured 3D model-image-text sequence, enabling the LMM to leverage past experiences and insights gained during previous iterations.
@@
In this representation, $P_{\rm iter}^{*},D_{\rm iter}^{*},F_{\rm iter}$ represent the optimal text prompt, 3D model, and textual feedback from each $\rm iter^{th}$ iteration. The memory module $m$ aids the agent in identifying specific T-2-3D output traits, such as misunderstood keywords. This knowledge is then integrated into generating refined 3D model prompts, enhancing the precision and adaptability of model generation. If T-2-3D struggles with certain design aspects, $m$ guides subsequent iterations to optimize prompts more effectively, ensuring continuous improvement in Idea23D, thereby increasing its alignment with complex IDEAs.
Table 1: We conducted experiments on the Eval3DAIGC-198 dataset with the configuration of generating one image per prompt ($num_{img}=1$), three prompts per round ($num_{draft}=3$), and up to five iteration rounds ($max_{iters}=5$). We used GPT-4o OpenAI (2024) instead of GPT-4-Vision-Preview OpenAI (2023a) from earlier studies. The models used include FLUX.1-dev Labs (2024), SD-XL 1.0 with refinement, Hunyuan3D-1.0 Yang et al. (2024), InternVL2.5-78B Chen et al. (2024b), and LLaVA-CoT-11B Xu et al. (2024a).
Comparison stage: T-2-3D vs. Avg. Iter. CLIP $\uparrow$ ULIP-2 $\uparrow$
LMM T-2-I I-2-3D T-2-3D Idea23D T-2-3D Idea23D
GT prompt FLUX InstantMesh - 0.3152 - 0.3134 -
GPT-4o FLUX InstantMesh 1.49 0.3003 0.3078 0.2733 0.2917
InternVL2.5 FLUX InstantMesh 2.56 0.2896 0.3035 0.2379 0.2756
LLaVA-CoT FLUX InstantMesh 4.67 0.2783 0.2826 0.2108 0.2125
Text Only FLUX InstantMesh - 0.2745 - 0.2056 -
GT prompt FLUX Hunyuan3D - 0.3085 - 0.3057 -
GPT-4o FLUX Hunyuan3D 1.41 0.2943 0.3010 0.2675 0.2869
InternVL2.5 FLUX Hunyuan3D 2.68 0.2870 0.2970 0.2534 0.2783
LLaVA-CoT FLUX Hunyuan3D 4.48 0.2734 0.2768 0.2189 0.2273
Text Only FLUX Hunyuan3D - 0.2700 - 0.2115 -
GT prompt SDXL InstantMesh - 0.2972 - 0.2684 -
GPT-4o SDXL InstantMesh 2.02 0.2845 0.3001 0.2302 0.2599
InternVL2.5 SDXL InstantMesh 3.49 0.2810 0.2822 0.2196 0.2211
LLaVA-CoT SDXL InstantMesh 4.53 0.2707 0.2726 0.1961 0.1991
Text Only SDXL InstantMesh - 0.2680 - 0.1979 -
GT prompt SDXL Hunyuan3D - 0.2941 - 0.2479 -
GPT-4o SDXL Hunyuan3D 2.14 0.2782 0.2903 0.2118 0.2404
InternVL2.5 SDXL Hunyuan3D 3.94 0.2725 0.2828 0.2143 0.2198
LLaVA-CoT SDXL Hunyuan3D 4.32 0.2711 0.2735 0.1911 0.1956
Text Only SDXL Hunyuan3D - 0.2663 - 0.1924 -
@@ 4.1 Experimental Setup
In our early experiments, the LMMs used were GPT-4V OpenAI (2023a) and LLaVA Liu et al. (2023b). Fig. 3, Fig. 7, Fig. 9, Fig. 10, and Fig. 11 use GPT-4V as the LMM, DALLE OpenAI (2023c) as the T-2-I model, and zero123 Liu et al. (2023d) as the I-2-3D model. The results of the User Study, shown in Tab. 3, were obtained from these early experiments. With the emergence of new LMM, T-2-I, and I-2-3D models, we have also tested Idea23D with these new methods. Fig. 1 and Fig. 8 use GPT-4o as the LMM, FLUX as the T-2-I model, and InstantMesh as the I-2-3D model. The quantitative results of the ablation study, presented in Tab. 2, were also based on these new models. This further demonstrates that Idea23D exhibits excellent compatibility with the development of LMMs, T-2-I, and I-2-3D models.
@@
T-2-3D baseline. Our first baseline is caption-based. Since no former 3D AIGC methods can be used for IDEA input, we convert image inputs and 3D model inputs (after multiview rendering) into textual descriptions by captioning them with the LMM. In Fig. 3, Tab. 1 and Tab. 3, this baseline is called T-2-3D. All T-2-3D baselines only use LMM for generating captions and do not perform iterative refinement.
@@
Ours w/o iterative refinement. To demonstrate the impact of an iterative self-refinement design, we construct an Idea23D variant with only one iteration. Compared with caption-based baselines, this one features multiple prompt generation and best 3D model selection. We presented the results of this part in our early user study, see Sec. 4.4.
4.2 Evaluation Dataset
In our experiments, we found that there is a lack of methods to align Text, Image, and 3D in the evaluation of LMM’s 3D AIGC capabilities Liu et al. (2023g); Andriushchenko et al. (2024); Guo et al. (2024); Wu et al. (2023b); Zhang et al. (2023a). Therefore, following the evaluation practices of Parti Yu et al. (2022), we constructed a dataset for evaluating 3D AIGC tasks, called Eval3DAIGC-198, which involves 198 different multimodal interleaved inputs of IDEA, including examples shown in Fig. 1, Fig. 3 and Fig. 8. The distribution and examples of the dataset can be found in Appendix A. The cases in the evaluation dataset consist of a text prompt, which may include images or 3D models. Since the results of 3D AIGC are difficult to represent by a specific 3D model as a standard answer, we annotated a description of the 3D model to be generated for each case based on the textual instructions, which serves as the Ground Truth.
4.3 3D-Caption Quantitative Results
To ensure a fair and comprehensive comparison of the alignment between user inputs and the final generated 3D models, we employed various methods to convert the generated 3D models into textual descriptions. These textual features were then compared with the handwritten 3D annotations in the Eval3DAIGC-198 evaluation dataset.
(1) CLIP Radford et al. (2021) Metric. We render the 3D models from four views (front, back, left, right) and calculate the CLIP similarity between the text description and the rendered images. The CLIP similarity is:
$S_{\text{CLIP}}(T,I)=\frac{1}{4}\sum_{i=1}^{4}\frac{E_{T}\cdot E_{I_{i}}}{\|E_{T}\|\|E_{I_{i}}\|}$ (9)
[…]
(2) ULIP-2 Xue et al. (2024) Metric. ULIP-2 is a tri-modal pre-training framework that generates text descriptions for 3D shapes without human annotations. It evaluates the alignment between 3D models and texts by comparing 3D shape features with corresponding descriptions in the Eval3DAIGC-198 dataset.
The quantitative experimental results are shown in Tab. 1. The "GT prompt" row represents the scores from generating 3D models using manually annotated 3D captions with text-to-image and image-to-3D methods. "Text Only" refers to using only textual instructions from the dataset as the baseline. The "T-2-3D" column shows results where LMM generates descriptions of images and 3D models, which are then combined with dataset instructions to generate 3D models. "Idea23D" represents our proposed method, and "Avg. Iter." shows the average number of optimization iterations for Idea23D.
The results demonstrate that Idea23D improves success rate and accuracy in 3D AIGC tasks, producing outputs closer to ground truth (GT prompt) with fewer iterations. Using GPT-4o as LMM, only two iterations are needed to generate realistic 3D models. Fig. 7 compares Idea23D with current commercial models, showing that Idea23D achieves new 3D AIGC capabilities not possible with existing models.
4.4 User Study
@@
We presented the results of the caption-based T-2-3D baseline and Idea23D to the participants in our user study. Our evaluation results are detailed in Tab. 3, which compares the caption-based T-2-3D baseline with Idea23D using the model specified in each row. We asked users to assess which model (T-2-3D model, first-round results of Idea23D, and final-round results of Idea23D) was more satisfactory, and to evaluate whether each 3D model complied with the IDEA. Detailed User Study results and explanations can be found in Appendix C.
4.5 Visualization of Self-iterative Refinement
Fig. 9 showcase the evolution and refinement of 3D models at different Idea23D stages and how the caption-based baseline works. Fig. 10 shows the iterative self-improvement process of a case in Idea23D. These visual representations illustrate the framework’s effectiveness in aligning with user IDEAs and the incremental improvements achieved through its iterative process.
4.6 Ablation Study
We conducted ablation experiments using the first 38 cases from the dataset, with LMM using GPT-4o, T2I using FLUX, and I23D using InstantMesh Iteration rounds and configurations are consistent with Tab. 1. Results are shown in Tab. 2. The case in Fig. 4 uses GPT-4V, SD-XL, and zero123.
As shown in the Pokemon case (Fig. 4), removing key modules slows convergence, while the full Idea23D model generates a satisfactory result within 3 iterations. The quantitative study in Tab. 2 reveals: (1) The memory module prevents quality divergence, (2) The LMM feedback agent accelerates convergence, (3) Removing stored information from previous models lowers the quality limit after convergence. Additionally, ablation results show that removing the memory module leads to increasing deviation from user inputs during iterations. Feedback improves convergence speed, and retaining prior models accelerates the process. The ULIP score for 3D models generated with real prompts is 0.3213, with Idea23D reaching 0.3208 after five iterations, while a standard text-to-image pipeline achieves only 0.2830.
Figure 4: Key module ablation across iterations. Note that the full model reaches satisfactory results (judged by the LMM agent 5 in Fig. 2) at iteration 3. The illustrated experiment has no maximum iteration limit.
Table 2: Ablation study results with the same configuration as in Tab. 1.
[…]
Iter 1 Iter 2 Iter 3 Iter 4 Iter 5
w/o memory 1.58 0.2953 0.2216 0.1902 0.2139 0.1960
w/o feedback 1.87 0.2930 0.2780 0.2759 0.2773 0.2811
w/o prev. model 1.78 0.2977 0.3031 0.3066 0.3012 0.2974
full model 1.49 0.3021 0.3056 0.3108 0.3146 0.3208
T-2-3D - 0.2830 - - - -
GT prompt - 0.3213 - - - -
Text Only - 0.2717 - - - -
5 Conclusion
Idea23D, utilizing an LMM agent collaboration framework, revolutionizes 3D AIGC by automating the creation of models from high-level, interleaved multimodal user inputs (IDEAs). This innovative system excels in integrating text, images, and 3D models, underpinned by a unique iterative process that enhances model coherence and visual alignment with IDEAs. User studies underscore its superiority in satisfaction and comparative quality, marking Idea23D as a significant advancement in 3D AIGC and a benchmark for future design tools.
6 Limitations
Idea23D effectively improves the alignment between 3D generation models and user intent, but it still relies on LMM and I-2-3D models. As shown in Tab. 1, LMM significantly impacts the 3D generation results. However, commercial models like GPT-4V can already generate image prompts based on feedback very well. On the other hand, open-source LMMs like LLaVA still have significant shortcomings in image understanding capabilities.
According to our experiments, the main bottleneck of Idea23D at this stage lies in the Image-to-3D step. Even the most advanced Image-to-3D models can fail. In the worst-case scenario with a very low probability, the final model output when Idea23D reaches the maximum iteration may still not meet user requirements. Nonetheless, Idea23D can ensure that the final selected model is the most aligned with user input among all generated 3D models throughout the iterations.
7 Acknowledge
This work was supported by the Fundamental Research Funds for the Central Universities under Grant Number KG16336301 and the China Postdoctoral Science Foundation under Grant Number 2024M764093.
Appendix A Appendix: Evaluation Dataset
These IDEAs span a range of complexities: 9 were text-only, 57 featured text and image inputs, 68 included text and 3D model inputs, and 64 contained text, image, and 3D model inputs. Each test case was meticulously designed to represent real-world scenarios. The dataset cases are the same as the IDEAs in Fig. 1 , Fig. 3 and Fig. 8.
Figure 5: IDEA Content Distribution.
The dataset also includes a distribution of tags: 9 IDEAs contained 0 tags, 62 IDEAs had 1 tag, and 127 IDEAs included 2 tags, highlighting the diversity of annotation complexities.
Figure 6: Tag Distribution in IDEAs.
Appendix B Appendix: Visualization Results
Figure 7: Comparison with commercial T-2-3D models. Inputs are the same as Fig. 3. Case (x) corresponds to the case of Fig. 1.
Figure 8: Results using GPT-4o, FLUX, InstantMesh as Idea23D components, cases using cases from the dataset.
The results of the visualization are shown in the Fig. 1, Fig. 7 and Fig. 8. For Tab. 3, we evaluate three T-2-I mdoels: DeepFloyd IF Lab (2023), DALL·E OpenAI (2023b), and SD-XL Podell et al. (2023), five I-2-3D models: Zero123 Guo et al. (2023); Liu et al. (2023d), Wonder3D Long et al. (2023), TripoSR Tochilkin et al. (2024), InstantMesh Xu et al. (2024b) and LGM Tang et al. (2024), and two LMM agent options: GPT-4V OpenAI (2023d, d) and LLaVA Liu et al. (2024, 2023b).
Our framework has good compatibility. After our initial experiments, advanced image-to-3D generation models such as TripoSR Tochilkin et al. (2024), InstantMesh Xu et al. (2024b), and LGM Tang et al. (2024) have emerged. Tab. 1 demonstrates that using these state-of-the-art models as the base modules for our T-2-I and I-2-3D results in significant improvements in Idea23D generation.
Appendix C Appendix: User Study Results
Table 3: Results of the user study.
@@ Appendix C Appendix: User Study Results
Comparison stage: T-2-3D vs. Idea23D (Initial-round) Idea23D (Iterative self-refined)
@@
Satisfying IDEA
T-2-I I-2-3D LMM T-2-3D Idea23D T-2-3D Idea23D T-2-3D Idea23D T-2-3D Idea23D
SD-XL TripoSR GPT-4V 38.5% 61.5% 54.7% 80.2% 18.9% 81.1% - 96.2%
SD-XL InstantMesh GPT-4V 43.2% 56.8% 53.1% 82.9% 20.1% 79.9% - 96.4%
SD-XL LGM GPT-4V 41.3% 58.7% 56.2% 81.3% 12.5% 87.5% - 94.5%
DALL·E Zero123 GPT-4V 25.8% 74.2% 41.5% 78.2% 6.5% 93.5% - 94.2%
DALL·E Zero123 LLaVA 27.5% 72.5% 33.2% 70.6% 29.7% 70.3% - 82.3%
DALL·E Wonder3D GPT-4V 18.8% 81.2% 47.9% 74.3% 3.5% 96.5% - 91.1%
DALL·E Wonder3D LLaVA 16.3% 83.7% 28.3% 69.4% 10.3% 89.7% - 76.2%
DeepFloyd IF Zero123 GPT-4V 20.0% 80.0% 38.5% 64.1% 28.3% 71.7% - 73.7%
DeepFloyd IF Zero123 LLaVA 29.0% 71.0% 23.1% 57.6% 34.0% 66.0% - 66.8%
DeepFloyd IF Wonder3D GPT-4V 29.3% 70.7% 32.6% 66.1% 14.8% 85.2% - 75.9%
DeepFloyd IF Wonder3D LLaVA 26.0% 74.0% 19.3% 55.3% 18.3% 81.7% - 64.5%
Tab. 3 shows the results of our user preference study. Our user study was conducted by distributing an online web-based survey questionnaire, with over 200 users participating. This user study’s comparative analysis reveals the remarkable superiority of the Idea23D framework over existing caption-based T-2-3D methods. Due to the lack of multi-round iterations in the T-2-3D baseline, the empty parts in the "Iterative self-refined" column of the table are the same as those in the "Initial-round" column.
Participants were presented with various IDEAs alongside multiple 3D model outputs from both caption-based T-2-3D baselines and Idea23D. To aid decision-making, users viewed rotating video representations of each 3D model. The order of presentation for both the cases and the model outputs was randomized to avoid any sequence bias, ensuring an unbiased assessment of the users’ preferences.
Our evaluation, detailed in Tab. 3, compares caption-based T-2-3D baselines and Idea23D using models specified in each row. We ask users to evaluate which model (the T-2-3D model, the first round results of Idea23D, and the end round results of Idea23D) is more satifying, as well as evaluating each 3D model for compliance with IDEA.
@@
The results show that Idea23D markedly enhances user preference scores across a diverse range of T-2-I, I-2-3D, and LMM models. Notably, the Idea23D framework’s initial prompting stage significantly outperforms the caption-based T-2-3D results by effectively decomposing and interpreting the user’s multimodal IDEA, thereby selecting the most suitable 3D model. This improvement is further amplified in the iterative self-refinement stage of Idea23D.
@@
For instance, in scenarios utilizing DALL-E, Zero123, and GPT-4V, Idea23D models were preferred by users in 74.2% of the cases over T-2-3D in the initial round, demonstrating a higher IDEA satisfaction rate (78.2%). Conversely, T-2-3D models achieved only 41.5% satisfaction. In the iterative refinement comparisons, Idea23D models were favored even more (93.5% preference), achieving a remarkable 94.2% satisfaction rate.
@@
These findings were consistent across various T-2-I, I-2-3D, and LMM configurations. Additionally, we observed that stronger T-2-I models, with enhanced language understanding capabilities, contributed to improved performance in Idea23D, suggesting that our framework may enjoy the development of off-the-shelf models it invokes.
Appendix D Appendix: Visualization of Self-iterative Refinement